OpenStudy (something):

How to solve? (g+6)/(6f^(3))÷(4g+24)/(3f)

OpenStudy (anonymous):

\[((g+6)/(6f{}^{\wedge}(3)))/((4g+24)/(3f))\text{ =}\frac{6+g}{2 f^2 (24+4 g)}\text{=}\frac{1}{8 f^2}\]